if z=-1 then the principal value of the Arg(z^2/3) is equal to

Grade:12th pass

1 Answers

Arun
6 years ago
z= –1
we can write z = cos(pi) + i sin(pi) = ei\pi
now z2/3 = ei 2\pi/3
hence arg(z2/3) = 2\pi/3
hope it helps